# APLe: Token-Wise Adaptive for Multi-Modal Prompt Learning

### 12 Jan 2024 v1

@@ 4 Experiments
To evaluate the capability of APLe, we conduct experiments with the well-known methods CLIP (30, 44, 43) and one recently proposed method multi-modal (17). The experiment and the datasets used align with prior research including three experiments and datasets of Caltech101 Caltech101 (7), DTD DTD (4), EuroSAT EuroSAT (11), FGVCAircraft FGVCAircraft (27), Flowers102 Flowers102 (28), Food101 Food101 (2), ImageNet ImageNet (5), OxfordPets OxfordPets (29), StanfordCars StanfordCars (20), SUN397 SUN397 (38), UCF101 UCF101 (33), ImageNetSketch object-detection4 (35), ImageNetV2 ImageNetV2 (32), ImageNet-A ImageNet-A (12) and ImageNet-R ImageNet-R (13).
@@ 4.1 Fundamental Experiment
Base-to-Novel Generalization: To examine the zero-shot learning capability, the datasets are split into two groups, base classes and novel classes. APLe is trained on the base classes in a few-shot manner and then evaluated in both groups.
Cross-Dataset Generalization: For the evaluation of zero-shot learning and transfer learning, we train APLe in ImageNet datasets by training all 1000 classes in a few-shot manner. Then, the evaluation is conducted in another 10 datasets.
Domain Generalization: To further demonstrate the performance of APLe on generalization capability, The model trained in cross-dataset generalization is evaluated for the other four datasets with a domain shift from ImageNet.
@@ 4.2 Base-to-Novel Generalization
Table.1 demonstrates the capability of APLe by comparison among existing prompt learning approaches including original CLIP with zero-shot CLIP (30), CoOp COOP (44), Co-CoOp COCOOP (43) and MaPLe multi-modal (17) over 11 datasets. We evaluate APLe by comparing to Co-CoOp and MaPLe separately to address the diverse characteristics of our approach among the other approaches.
Base Classes in Generalization: While performing prompt learning, Co-CoOp drops the average accuracy of 11 datasets in base classes from 82.69% to 80.47% compared to CoOp. By the independent and sequential framework, APLe stabilized the average accuracy at 81.99% with an advanced generalization capability in all datasets compared to Co-CoOp which demonstrates the capability in line with the baseline.
Novel Classes in Generalization: In novel classes, While APLe applied conditioning effort on the image branch and incorporated zero-shot knowledge, APLe indicated significant improvement over he datasets except for the OxfordPets with the 0.13% drop in novel class. Over 11 datasets, the average accuracy of APLe in novel classes generalization is boosted from 71.69% to 75.11%. Meanwhile, APLe demonstrates a 2.57% significant increase if both base classes and novel classes are considered which is denoted as harmonic means.
@@
Table 2: Effectiveness analysis of sigma in two failure cases.
APLe $vs$ MaPLe: In Table.1 (a), it indicates that the average accuracy is comparable between APLe and MaPLe, MaPLe applied multi-modal prompting with a coupling function. APLe outperforms the advantage in 5/11 datasets and has equivalent performance in the other 4 datasets (average gap = $\sim$0.32%) on average accuracy. The disadvantage occurs at DTD and FGVCAircraft with 1.15% and 1.16% gap respectively. In general, APLe shows competitive performance in 4/11 datasets and comparable results in the other 5 datasets (average gap $\sim$ 0.20%). In DTD and FGVCAircraft, the drop happened with 0.97% and 0.98% respectively.
@@
Failure Cases: We analyze cases where the APLe fails while MaPLe succeeds. Specifically, Taking into account the non-learnable nature and sensitivity to image style, we assess the effectiveness of different sigma used in the image adapter as shown in Table. 2. The table shows that APLe achieves advanced performance when sigma is set to 0.02 for DTD and 0.12 for FGVCAircraft. Considering the nature of the Gaussian filter used in the image adapter, the results suggest that images from DTD and FGVCAircraft are situated at two distinct extremes. One necessitates preserving high-frequency details and edge features, while the other calls for more extensive smoothing to alleviate noise. As the sigma decreases, APLe enhances performance in accuracy for both base and novel classes in the DTD dataset while the increase in sigma improves the generalization capability in FGVCAircraft. In section 4.5, the further characteristics and advantages of APLe are explored in detail to adopt V-L models.
@@ 4.3 Cross-Dataset Generalization
We evaluate the generalization performance between datasets from ImageNet to other 10 datasets by training 1000 classes. Table 3 demonstrates the ability among CoOp, Co-CoOp and MaPLe. APLe exhibits a modest result on the source dataset, yet ultimately demonstrated a commonly stronger competitive performance of 66.23% with 6/10 datasets surpassing Co-CoOp and 5/10 datasets surpassing MaPLe. APLe achieves equivalent performance in other datasets and outperforms 3 top records. This indicates that the APLe achieves comparable generalization performance with other methods.
@@ 4.4 Domain Generlization
By training the model on the ImageNet and evaluating the modal in the out-of-domain datasets, Our approach excels in the generalization performance on out-of-distribution datasets among CoOp, Co-CoOp and MaPLe. The average results of APLe are among the top as indicated in Table 4 argues that deploying the token-wise knowledge training with adaptation for multi-modal prompt supports APLe to attain a competitive generalization capability among the methods.
@@ 4.5 Vital Experiments and In-Depth Analysis
Prompt Length and Overfitting: We initially explored the performance of vision and language as well as both prompts across different prompt lengths shown in Table. 5 with a simplified structure demonstrated As the prompt length increases for both prompts or solely for the vision prompt, the generalization performances experience minor fluctuations along with a modest decrease. Meanwhile, the increase in text prompt length leads to a noticeable decrease in accuracy. In other words, the increase in prompt length in the language prompt has a greater impact compared to that of the vision prompt while these effects are often characterized by a decrease in accuracy. The impact of decreasing performance is commonly described as overfitting in existing papers multi-modal (17) due to prompt conditions such as length. The results suggest that prompt learning behavior varies across the two different modalities including generalization capability and the risk of knowledge conflicts. Owing to the designed framework, APLe effectively improves the generalization performance and mitigates the adverse influences of challenges from the divergence of modalities token and the overfitting introduced by prompt length grows in training while both challenges are well discussed in existing researches.
We evaluate the generalization capability of APLe over different prompt lengths and introduce two evaluations for generalization evaluation. The harmonic mean indicates the generalization performance as foundational experiments while the standard deviation effectively measures the variability in predictive results across 7 different prompt lengths. The evaluations provide an assessment of overfitting for the models. In Table.6, the average performance of APLe provides significant advantages over MaPLe in both accuracy and standard deviation. The average accuracy over 7 different lengths and its standard deviation to the population are 77.07% and 1.37 for APLe, 76.68% and 1.92 for MaPLe. In Fig. 3, the equivalent outcomes repeat when illustrating the generalization performance in two chosen datasets containing numerous textures and small objects and containing the most categories and fine-grained, EuroSAT and StanfordCars, where APLe indicates the more stable curves in both evaluations while the detailed evaluation is demonstrated. The outcomes show that APLe’s framework assists it in further improving generalization capability over the baselines.
@@
Figure 4: noAdapt $vs$ Adapt Comparison.
Effectiveness of Multi-modal Token Adaptation: The designed framework incorporates CLIP zero-shot knowledge to encourage APLe to learn the knowledge comprehensively and independently while alleviating overfitting challenges caused by prompt settings. However, this potentially leads to difficulties in effective collaboration between two tokens. Thus, we introduce token adaptation to promote the synergy between prompt learning and further enhance the model’s generalization performance. In Fig. 4, the results of the analysis in chosen datasets, DTD, EuroSAT, OxfordPets and ImageNet, as well as average performance over 11 datasets, exhibit significant improvement in base classes and overall performance, which suggests the token adaptation supports APLe in improving the generalization capability during prompt learning and provide advanced performance in downstream tasks.
@@
Effectiveness of hyper-parameter $\lambda$: We further analyze the effect of the hyper-parameter $\lambda$ ($d$ and $g$) used in cooperating zero-shot knowledge for corresponding modality, language and vision respectively, token training as described in Eq. 1819. Results are shown in Table. 7. As pointed out previously, various feature representation complexities exist between the vision and language modalities. The distinctions are demonstrated explicitly in generalization performance and the risk of knowledge conflicts in extended experiments in prompt length. For the results shown in Table. 7, the accuracy in DTD, another dataset containing textures than EuroSAT, improves with increasing values of $\lambda_{d}$ while the result experiences fluctuations when $\lambda_{g}$ also rise. Alternatively, when $\lambda_{d}$ is in the vicinity of 0.8, accuracy achieves higher values. For OxfordPets with the fewest categories and fine-grained, the result indicates a similar pattern to achieve higher value, while the $\lambda_{d}$ is around 0.6. The result suggests that the prompt in the language modality shows a more stable generalization capability for knowledge learning and knowledge conflict resolution than the prompt in the vision modality.
@@ 5 Conclusion
The fundamental concern of prompt learning lies in the sensitivity to the setting, such as prompt lengths, while the extended prompts carry the risk of model overfitting. Likewise, the divergences between modalities contribute to the challenge of learning, which also further heightens the risk of overfitting. To this end, APLe introduced the sequential token-wise knowledge training and adaptation framework to address the challenges of multi-modal prompt learning. The proposed approach achieves comparable performance compared to the current state-of-the-art methods in three fundamental tasks. Particularly in datasets with domain shifts, indicates a significant gain in average accuracy. Of greater importance, the results in the analysis of the prompt length and overfitting exhibit APLe dominate the generalization performance across different prompt lengths in downstream tasks.

### 23 Jan 2024 v2

@@ 4 Experiments
To evaluate the capability of APLe, we deploy experiments with the aforementioned approaches, including CLIP, CoOp, Co-CoOp, and MaPLe. The experiment and the datasets used align with the baseline including three experiments and 15 datasets of Caltech101 Caltech101 (6), DTD DTD (3), EuroSAT EuroSAT (10), FGVCAircraft FGVCAircraft (26), Flowers102 Flowers102 (27), Food101 Food101 (1), ImageNet ImageNet (4), OxfordPets OxfordPets (28), StanfordCars StanfordCars (19), SUN397 SUN397 (38), and UCF101 UCF101 (35); ImageNetSketch object-detection4 (37), ImageNetV2 ImageNetV2 (31), ImageNet-A ImageNet-A (11) and ImageNet-R ImageNet-R (12) with domain shift from the ImageNet dataset.
@@ 4.1 Fundamental Experiment
To examine the zero-shot learning capability, the datasets are split into two groups, base classes and novel classes, with the evaluation of the average accuracy over 3 runs, and the harmonic mean (HM) of two accuracies is also included.
Base-to-Novel Generalization. APLe is trained on the base classes with 16-shot and then evaluated in both groups.
Cross-Dataset Generalization. For the evaluation of zero-shot learning and transfer learning, we train APLe in ImageNet by deploying all 1000 classes in a 16-shot manner. Then, the evaluation is conducted in another 10 datasets.
Domain Generalization. To further demonstrate the performance of APLe on generalization capability, The model trained in cross-dataset generalization is evaluated in the other four datasets with domain shift from the ImageNet.
@@ 4.2 Base-to-Novel Generalization
Table.1 indicates the capability of APLe, highlighted in grey, with original CLIP, CoOp, Co-CoOp, and MaPLe over 11 datasets. We compare APLe with Co-CoOp and MaPLe separately in performance to address the diverse features of APLe.
Base Classes in Generalization. While performing prompt learning, Co-CoOp drops the average accuracy of 11 datasets in base classes from 82.69% to 80.47% compared to CoOp. By the independent and sequential framework, APLe stabilized the average accuracy at 81.99% with an advanced generalization capability in all datasets compared to Co-CoOp, which demonstrates the capability in line with the others.
Novel Classes in Generalization. Compared to Co-CoOp in novel classes, owing to applying conditioning effort on the image branch and incorporating zero-shot knowledge, APLe shows significant improvement except for the OxfordPets with the 0.13% drop in the novel class. Over 11 datasets, the average accuracy of APLe in novel classes generalization is boosted from 71.69% to 75.11%. Meanwhile, APLe demonstrates a 2.57% significant increase if both base classes and novel classes are considered which is denoted as HM.
@@
Table 2: Effectiveness analysis of sigma $\sigma$ for two failure cases. The result highlighted denotes the values used in our case.
APLe $vs$ MaPLe. In Table. 1 (a), it indicates that the average accuracy is comparable between APLe and MaPLe, while both approaches apply multi-modal prompt learning. APLe demonstrates the advantage in 5/11 datasets and has equivalent performance in the other 4 datasets (average gap = $\sim$0.32%) in the novel classes. The disadvantage occurs at DTD and FGVCAircraft with 1.15% and 1.16% gap respectively. In general, APLe shows competitive performance in 4/11 datasets and comparable results in the other 5 datasets (average gap = $\sim$0.20%). In DTD and FGVCAircraft, the disadvantages amounted to 0.97% and 0.98% respectively.
@@
Failure Cases. We analyze the cases where the APLe fails while MaPLe succeeds. Specifically, Taking into account the non-learnable nature and sensitivity to image style, we assess the effectiveness of different sigma used in the Gaussian filter as shown in Table. 2. The table shows that APLe enhances performance in accuracy for both base and novel classes in the DTD dataset as the sigma decreases, while the increase in sigma improves the generalization capability in FGVCAircraft. APLe achieves advanced performance when sigma is set to 0.02 for DTD and 0.12 for FGVCAircraft. Considering the nature of the Gaussian filter used in the image adapter, the results suggest that images from DTD and FGVCAircraft are situated at two distinct extremes. One necessitates preserving high-frequency details and edge features, while the other calls for more extensive smoothing to alleviate noise. Besides the analysis of sigma in image adapter, the further characteristics and advantages of APLe are explored in detail in section 4.5.
@@ 4.3 Cross-Dataset Generalization
We evaluate the generalization performance between datasets from ImageNet to other 10 datasets by training 1000 classes. Table. 3 demonstrates the ability among CoOp, Co-CoOp and MaPLe. APLe exhibits a modest result on the source dataset and demonstrated a competitive generalization performance in 10 datasets by surpassing Co-CoOp in 6 datasets and surpassing MaPLe in 5 datasets. Ultimately, APLe outperforms 3 top records with an average accuracy of 66.23%. This indicates that the APLe achieves comparable generalization performance with other methods in cross-dataset evaluation.
@@ 4.4 Domain Generlization
By training the model on the ImageNet and evaluating the modal in the out-of-domain datasets, Our approach excels in the generalization performance on out-of-distribution datasets among CoOp, Co-CoOp and MaPLe. The average results of APLe are among the top as indicated in Table. 4. Owing to deploying the token-wise knowledge training with adaptation for multi-modal prompts, APLe attains a competitive generalization capability among the methods.
@@ 4.5 Vital Experiments and In-Depth Analysis
Prompt Length and Overfitting. We initially explored the performance of vision and language as well as both prompts across different prompt lengths shown in Table. 5 with a simple structure without token-wise training and adaptation. As the prompt length increases for both prompts or solely for the vision prompt, the generalization performances experience minor fluctuations along with a modest decrease. Meanwhile, the increase in text prompt length with leads to a noticeable decrease in accuracy. In other words, In the language prompt, extending the prompt length has a more pronounced effect, with nearly achieving the lowest scores highlighted in blue at each length. The effects are often characterized by a decrease in accuracy. The impact of decreasing performance is commonly described as overfitting in existing papers due to prompt conditions such as length. The results suggest that prompt learning behavior varies across the two different modalities, including generalization capability and the risk of knowledge conflicts. Owing to the designed framework of sequential token-wise knowledge training and token adaptation, APLe effectively improves the generalization performance and mitigates the adverse influences of challenges in the divergence of modalities token and the overfitting introduced by an increasing prompt length to adopt the V-L model.
We evaluate the generalization capability of APLe with MaPLe over different prompt lengths by two evaluation metrics. The harmonic mean indicates the generalization performance as foundational experiments, while the standard deviation effectively measures the variability in predictive results across 7 different prompt lengths. The evaluations provide an assessment of overfitting for the models. In Table. 6, the average performance of APLe provides significant advantages over MaPLe in both accuracy and standard deviation. The average accuracy over 7 different lengths and its standard deviation to the population are 77.07% and 1.37 for APLe, 76.68% and 1.92 for MaPLe. Specifically in Fig. 3, APLe exhibits notably stable curves for the HM in EuroSAT and StanfordCars. One contains numerous textures and small objects, and the other comprises the most categories with fine-grained.
In the exploration of the overfitting experiment, the results resoundingly attest to the effectiveness of the framework in empowering APLe to enhance its generalization capability, showcasing outstanding robustness compared to the baseline.
Effectiveness of Multi-Modal Token Adaptation. The designed framework incorporates CLIP zero-shot knowledge to encourage APLe to learn the knowledge comprehensively and independently while alleviating overfitting challenges caused by prompt settings. However, this potentially leads to difficulties in effective collaboration between two modalities. Thus, we introduce token adaptation to promote the synergy between prompts and further enhance the generalization performance. In Fig. 4, the results of the analysis in chosen datasets, DTD, EuroSAT, OxfordPets, and ImageNet, as well as average performance over 11 datasets, exhibit significant improvement in base classes and overall performance, which suggests the token adaptation supports APLe in improving the generalization capability during prompt learning and provide advanced generalization performance in downstream tasks.
@@
Figure 4: Adaptation $vs$ Non-Adaptation Comparison.
@@
Effectiveness of Hyper-Parameter $\lambda$. We further analyze the effect of the hyper-parameter $\lambda$ ($d$ and $g$) used in integrating the result from token-wise training, language and vision respectively, with zero-shot knowledge. as described in Eq. 1819, while results are shown in Table. 7. As pointed out previously, various feature representation complexities exist between the vision and language modalities, the distinctions are shown explicitly in performance and the risk of knowledge conflicts in extended experiments in prompt length. For the results shown in Table. 7, the accuracy in DTD, another dataset containing textures than EuroSAT, increases with increasing values of $\lambda_{d}$, but the result is expected to experience fluctuations if $\lambda_{g}$ rise. Alternatively, the accuracy achieves higher values when $\lambda_{d}$ is in the vicinity of 0.8, highlighted in blue. For OxfordPets with the fewest categories with fine-grained, the result shows a similar pattern, it achieves a higher value while the $\lambda_{d}$ is around 0.6. The results suggest that the prompt in the language modality presents a more stable generalization capability for knowledge learning and knowledge conflict resolution than the prompt in vision.
@@ 5 Conclusion
The fundamental concern of prompt learning lies in the sensitivity to the setting, such as prompt lengths, while the extended prompts carry the risk of model overfitting. Likewise, the divergences between modalities contribute to the challenge of learning, which also further heightens the risk of overfitting. To this end, APLe introduced the sequential token-wise knowledge training and adaptation framework to address the challenges of multi-modal prompt learning. The proposed approach achieves comparable performance compared to the current state-of-the-art methods in three fundamental tasks. Particularly in datasets with domain shifts, the result indicates a significant gain in average accuracy. Of greater importance, the results in the analysis of the prompt length and overfitting exhibit APLe dominates the generalization performance across different prompt lengths. In light of the preceding experiments, APLe exhibits substantial generalization capability and emerges as a robust and adaptable approach for adopting the V-L model in downstream tasks.